# LLM-based Triplet Extraction for Automated Ontology Generation in Software Engineering Standards

### v1

@@ 4.1. Systems Compared
Pred-LLM is the result from this research–Figure 2(e) panel (d). Open-source 7B LLM (Mistral-7B-Instruct-v0.1 (Jiang et al., 2023)) is used for open-set relation generation.
@@
Pred-openIE is a sentence-level OpenIE baseline–an example is shown in Figure 2(e) panel (e). Stanford OpenIE (Angeli et al., 2015) implemented in Stanford CoreNLP (Manning et al., 2014) is employed.
4.2. Evaluation Protocol
@@
Fully automated outputs (no post-editing) are evaluated against three gold-standard reference sets for the short SECEPP: Gold-Short, Gold-Medium, Gold-Long (Figure  2(e); stats in Table 1).
@@
Quality is assessed at the node and triple levels via an embedding-similarity threshold sweep $\tau\in\{0.10,0.15,\dots,0.90\}$; Precision, Recall and F1 are reported as curves (Fig. 3, top: nodes; bottom: triples; columns: Short/Medium/Long). Matches require the similarity of the concatenation of subject, relation, and object (a single space delimiteris used) to pass $\tau$ (greedy one-to-one alignment).
4.3. Key Findings
@@
(1) Recall is highest on Gold-Long and lowest on Gold-Short, confirming the recall-stress hypothesis for dense gold sets. The Gold-Long gold is a compact backbone, and it eases recall (fewer gold edges), but any extra predicted edges not in the backbone are counted as false positives, lowering precision.
@@
(2) Pred-LLM often yields higher precision than OpenIE across $\tau$, and only when $\tau\!>\!0.5$, the recall gap narrows, which makes sense as Pred-LLM’s result has a much smaller number than Pred-openIE’s result ($42$ vs. $74$ under Gold-Short) and spurious/duplicate triples from OpenIE are filtered.
(3) It is observed that starting from $\tau\approx$0.6, the matching of the triples is highly trustworthy. This is an interesting finding. Further work will investigate the optimal $\tau$ value as the most effective indicator for this type of study.
@@
(5 )A 7B open-source LLM is already competitive with OpenIE—higher precision on both nodes and triples, and comparable recall on stricter gold sets (when $\tau\!>\!0.5$)—while recall on Gold-Short remains the main headroom to be improved.
@@
(6) Triple matching is strictly harder than node matching: errors in any of head/relation/tail cause a miss, so generally both precision and recall drop at the triple level (error compounding).
5. Conclusion / Future Work
@@
This work initiates LLM-assisted ontology generation for software engineering standards, proposes a fully automated workflow, constructs three gold-standard reference sets for the short version, and evaluates with threshold-sweep metrics at node and triple levels. Preliminary results show that a 7B open-source LLM is competitive with a Stanford OpenIE baseline (often higher precision, comparable recall on stricter gold sets).

### v2

@@ IV-A Systems Compared
Pred-LLM is the result from this research–Figure 2 panel (d) is one representative output out of five runs, the mean performance of which is reported in Figure 3. An open-source 7B LLM (Mistral-7B-Instruct-v0.1 [10]) is used in this study of open-set relation generation. Code and data are available at https://github.com/songhui01/aog-ses.git.
@@
Pred-openIE is a sentence-level OpenIE baseline–an example is shown in Figure 2 panel (e). Stanford OpenIE [3] implemented in Stanford CoreNLP [19] is employed.
IV-B Evaluation Protocol
@@
Fully automated outputs (no post-editing) are evaluated against three ref-standard reference sets for the short SECEPP: Ref-Short, Ref-Medium, Ref-Long (Figure  2; stats in Table I).
@@
Quality is assessed at the node and triple levels via an embedding-similarity threshold sweep $\tau\in\{0.10,0.15,\dots,0.90\}$; Precision, Recall and F1 are reported as curves (Figure 3, top: nodes; bottom: triples; columns: Short/Medium/Long). Matches require the similarity of the concatenation of subject, relation, and object (a single space delimiteris used) to pass $\tau$ (greedy one-to-one alignment).
IV-C Key Findings
@@
(1) Recall is highest on Ref-Long and lowest on Ref-Short, confirming the recall-stress hypothesis for dense reference sets. The Ref-Long is a compact backbone, and it eases recall (fewer edges), but any extra predicted edges not in the backbone are counted as false positives, lowering precision.
@@
(2) Pred-LLM often yields higher precision than OpenIE across $\tau$, and only when $\tau\!>\!0.5$, the recall gap narrows, which makes sense as Pred-LLM’s result has a much smaller number than Pred-openIE’s result ($42$ vs. $74$ under Ref-Short) and spurious/duplicate triples from OpenIE are filtered.
(3) An additional observation from the experimental checks is that when $\tau$ reaches around 0.6, the triples generally stabilize and become reliable. The supporting evidence could not be included in this short paper, but this preliminary finding motivates further investigation of the optimal $\tau$ value as the key indicator for this type of study.
@@
(4 )A 7B open-source LLM is already competitive with OpenIE—higher precision on both nodes and triples, and comparable recall on stricter reference sets (when $\tau\!>\!0.5$)—while recall on Ref-Short remains the main headroom to be improved.
@@
(5) Triple matching is strictly harder than node matching: errors in any of head/relation/tail cause a miss, so generally both precision and recall drop at the triple level.
IV-D Discussion
The results suggest that a 7B open-source LLM can already be competitive with OpenIE in precision for node and triple extraction. This indicates that stronger LLMs—either with larger parameter sizes or fine-tuned on domain-specific data—may further improve extraction quality.
However, this preliminary study has several limitations. First, it evaluates the workflow using a single LLM and thus lacks comparative experiments across LLM families and scales, as well as across backbone variants such as dense vs. MoE. Second, the reference datasets are manually constructed by a single annotator (domain expert) at three different granularities to support systematic evaluation in the absence of existing gold sets for this domain/task; consequently, inter-annotator agreement (IAA) is not reported. These labels are intended for formative evaluation rather than a community benchmark. In future work, we will include multiple annotators and report IAA to improve label reliability and support the construction of a public benchmark dataset.
V Conclusion / Future Work
@@
This work initiates LLM-assisted ontology generation for software engineering standards, proposes a fully automated workflow, constructs three expert-annotated reference sets (single annotator in this priliminary study) for the short version, and evaluates with threshold-sweep metrics at node and triple levels. Preliminary results show that a 7B open-source LLM is competitive with a Stanford OpenIE baseline (often higher precision, comparable recall on stricter reference sets).
@@
VI Acknowledgements
This material is based upon work supported by the National Science Foundation under Grant No. 2451397.
AI assistants (OpenAI’s GPT-4.x and Google’s Gemini suggestions embedded in Google Colab) were used for rapid prototyping during the experiments. No AI was used to fabricate data or results. The code was comprehensively reviewed and revised by the author. AI tools were also used to rephrase some passages for conciseness and grammatical correctness.